What is the radical form of 8/3/5 (8 and 3 fifths)
I was told that it is 24/5 but I can't understand the steps to the answer

Answer by Alan3354(69443)

8 3/5 = 8 + 3/5
8 = 40/5
--> 40/5 + 3/5
= 43/5
-------
24/5 = 2 and 4/5
There are no steps to get that.
I don't think this is called the "radical form", it's an "improper fraction."